OpenStudy (anonymous):

Simplify the expression. Assume that all variables are positive. Problem below:

OpenStudy (anonymous):

\[\sqrt[3]{x}^7 . \sqrt[3]{x}^2\]

OpenStudy (jit4won):

cube root of x^9 = x^3

OpenStudy (anonymous):

\[x ^{7/3}.x ^{2/3}=x ^{7/3 +2/3} =x ^{9/3} =x ^{3}\]